OpenStudy (anonymous):

solving systems using substitution x=y x+2y=3

OpenStudy (sandra):

hey julian

OpenStudy (anonymous):

hey sandra

OpenStudy (sandra):

so to solve by substitution, you replace one of the variables in a second equation with the value for that same variable in the first equation. This replacement is called substitution.

OpenStudy (sandra):

so in your example, we want to replace the x in the second equation, with the value of "x" in the first - i.e. subsitute for x in the second equation

OpenStudy (anonymous):

ok

OpenStudy (sandra):

so instead of x + 2y =3, we use "x=1" in the first equation to replace the "x" with "y" 1. x + 2y = 3, and given x=y 2. (y) + 2y = 3 3. 3y = 3 4. y = 1

OpenStudy (sandra):

in general, all substitution problems come down to "isolating" a variable in one of the given equations (getting it by itself - in this problem it's already started that way with x=y), and then using that value in place of the same variable in another equation. replacement = subsitution

OpenStudy (anonymous):

oh ok im staring to get it now

OpenStudy (anonymous):

Thank you!!!!! you really helped

OpenStudy (sandra):

ok np, post more problems if you want to work through another one!

OpenStudy (anonymous):

ok